# Suppression of errors in collectively coded information

### v1

@@ Abstract
Modern life largely transmits genetic information from mother to daughter through the duplication of single physically intact molecules that encode information. However, copying an extended molecule requires highly processive copying machinery and high fidelity that scales with the genome size to avoid the error catastrophe. Here, we explore these fidelity requirements in an alternative architecture, the virtual circular genome, in which no one physical molecule encodes the full genetic information. Instead, information is encoded and transmitted in a collective of overlapping and interacting segments. Using a model experimental system of a complex mixture of DNA oligos that can partly anneal and extend off each other, we find that mutant oligomers are suppressed relative to a model without collective encoding. Through simulations and theory, we show that this suppression of mutants can be explained by competition for productive binding partners. As a consequence, information can be propagated robustly in a virtual circular genome even if the mutation rate is above the error catastrophe for a physically intact genome.
Modern life transmits genetic information using long, physically continuous polymers such as chromosomes and plasmids. This mode of storage imposes several stringent demands on the copying process. First, replication must be highly processive: subsequent bases of the coding polymer must be copied in sequence without dissociation. Second, the strands must be physically separated during copying, requiring mechanisms capable of strand displacement. Third, fidelity must scale and increase with genome length to avoid the so-called error catastrophe, where errors accumulate faster than they can be purged by natural selection[1].
All of these requirements are met by complex life today but at the cost of enormous complexity in the copying machinery. For example, the requirements of strand displacement and processivity are are handled by the coordinated action of helicases, single-strand binding proteins and polymerases. Fidelity is enforced through multiple layers of error correction, from nucleotide selectivity[2] and exonucleolytic proofreading[3] to post-replicative mismatch repair[4]. But such complexity likely arose gradually and would have been unavailable during the earliest stages of evolution of life on earth, raising the question of how genomic information could be copied at early stages. Notably, even efforts to engineer minimal cells today are constrained by these same requirements: currently, these ‘minimal’ cells must still encode genomes large enough to code for complex proteins required for reliable replication of long polymers[5].
As a consequence, several prior studies have explored how informational patterns can emerge and persist in pools of short oligonucleotides, often using random sequence mixtures subjected to cycles of annealing, extension, or ligation to mimic simple or primitive replication environments[6, 7, 8, 9, 10, 11, 12]. These studies showed that even in the absence of long templates or sophisticated enzymes, simple chemical processes can give rise to heritable molecular features. Building on this foundation, we examine a more structured scenario in which the genome is deliberately encoded across a designed set of overlapping sequences that form a virtual circular genome (VCG)[13, 14]. In this architecture, no single molecule encodes the full sequence; instead, information is distributed redundantly across a population of short oligonucleotides, each of which can act as both primer and template. This compositional, collectively encoded system enables replication through thermally-driven melt-anneal-extend cycles.
Here, we explore the propagation of errors in this alternative distributed architecture, the Virtual Circular Genome, for storing and replicating genetic information. We show that cooperative interactions inherent to the VCG architecture naturally suppress the propagation of replication errors. Using an experimental DNA system combined with simulations and theoretical analysis, we demonstrate that mutant alleles, particularly when they arise at 3-ends of strands, are at a competitive disadvantage (compared to wildtype alleles) due to reduced access to productive extension partners. This fidelity arises from ecological interactions among molecules, rather than explicit proofreading, and is enhanced by stalling at mismatches[15, 16]. These results suggest that the architecture of genome encoding can play a central role in enabling robust information transmission, with implications for early life and the design of minimal or synthetic self-replicating systems.
@@
Figure 1: Co-operative replication of genetic information in a virtual circular genome. (A) (i) In physically circular genomes found in extant life, genetic information is encoded in long nucleic acid polymers (ii) that are replicated by the extension of short priming strands. (iii) As each primer is extended to cover the entire genome, a neutral mutant allele in one part of the genome is replicated just as often as its wildtype counterpart. (B) (i) In contrast, the proposed virtual circular genome of a protocell is the consensus sequence (purple) of many short oligonucleotides (black and grey). (ii) Each oligonucleotide (oligo) may act as both primer and template during replication. (iii) Here, we show that the resulting co-operative effects, in which wildtype and mutant oligos compete as both primers and templates, suppress the replication of isolated mutant alleles in favor of wildtype alleles already coded on many oligos.
@@ I Experimental model
Figure 2: Suppression of mutant propagation in an experimental DNA model of the virtual circular genome (VCG). (A) 60 bp consensus sequence for the VCG analyzed here; sequence designed to avoid repeats of length $\geq$ 4 bp. VCG was synthesized as 24 DNA oligonucleotides (12 each clockwise and counterclockwise), each one 25 bp in length and staggered 5 bp along the consensus sequence. (B) Mutant oligos were designed by replacing a 4 bp region of one VCG oligo with a mutated sequence. (C) Experimental setup for VCG replication. Mix of VCG and mutant oligos (here, initial mutant oligo concentration is 5% of corresponding wildtype (wt) concentration) is combined with dNTPs (1mM), Bst DNA polymerase, and 1X Bst buffer and subject to thermal cycling. (D) Denaturing gel electrophoresis of samples after 0, 10, 20, and 30 thermal cycles. DNA ladders (outermost lanes) range from 25 to 90 nt. The initial 25 nt oligos at cycle 0 progressively extend, with predominant products reaching $\sim 60$ nt by 20 cycles, consistent with the melting temperature $T_{\mathrm{m}}=80~^{\circ}\mathrm{C}$ used in the cycles. (E) Amplification of wildtype and mutant oligos as a function of thermal cycle, for 5% initial mutant levels. Replication is quantified by qPCR every 10 thermal cycles, using primers specific to either a wildtype (green) or mutant (orange) sequence. Measured cycle threshold (Ct) values are converted to absolute DNA concentrations by a standard curve obtained by serial dilution of a known concentration. Amplification of an oligo is defined as oligo concentration normalized to its initial concentration before thermal cycling. Here, as later, circle and square markers denote two independent thermal-cycle replicates, each measured by qPCR in duplicate.
Although our ultimate interest lies in RNA-based virtual circular genomes, we employ here a DNA-based experimental model as done in prior related work[9, 12, 17, 18]. This choice allows precise control over sequence composition, strand concentrations, and reaction conditions, while enabling direct quantification of wildtype and mutant abundances. Importantly, despite the chemical differences between DNA and RNA, the critical features relevant to mutation suppression are preserved: both systems involve complex mixtures of overlapping oligomers that anneal to multiple partners, and relax slowly toward equilibrium through numerous competing interactions (including strand displacement). As such, the DNA system can faithfully test whether collective encoding suppresses mutation propagation, while deferring RNA-specific differences, such as persistence length and the mechanism of template copying, to later discussion.
We designed a DNA-based system composed of short, overlapping oligonucleotides that form a redundant circular architecture. The full VCG comprises 12 distinct 25-nt strands and their reverse complements. Each of the 12 strands in one direction overlaps with its downstream neighbor (i.e., the next strand in the 5–3 direction) by 20 bases, encoding a 60 bp circular genome (Fig. 2A, Supp. Table I). Thus, each strand initially has four binding partners that productively allow it to be extended or part of it to be copied, i.e., reverse-complementary strands that can hybridize and serve as templates and/or as primers, with partial complementarity ranging from 5 to 20 bp. Each strand also has an additional five binding partners which anneal in unproductive configurations that do not allow for primer extension.
To evaluate how mutations propagate within this architecture, we introduced a 25-nt mutant strand that differs from one VCG strand by a single 4 bp substitution near the center of the sequence (Fig. 2B, Supp. Sec. I). Mutant strands were spiked into VCG mixtures at defined low proportions (2.5%, 5%, or 25% relative to the wildtype) to mimic rare variant emergence.
Extension of VCG strands was driven by thermal cycling with Bst DNA Polymerase (Large Fragment, NEB M0275), using 10–30 cycles of alternating denaturation ($80~^{\circ}\mathrm{C}$) and annealing/extension ($35~^{\circ}\mathrm{C}$), followed by a final enzyme deactivation step at $90~^{\circ}\mathrm{C}$ for 10 minutes. Notably, no new oligomers are supplied to this system; extension is initiated by overlap-driven annealing of partially complementary VCG strands (Supp. Sec. III). Gel electrophoresis confirmed that initial 25-nt VCG strands were extended incrementally over cycles, with products reaching a length of 45-nt by cycle 10, and further reaching  60-nt by cycle 20, with gel profiles remaining unchanged through cycle 30 (Fig. 2E). Duplexes up to  45 bp ($T_{\mathrm{m}}=80~^{\circ}\mathrm{C}$) likely remain partially meltable during cycling, allowing strand reshuffling and continued extension, whereas  60 bp duplexes ($T_{\mathrm{m}}=82~^{\circ}\mathrm{C}$) are too stable to denature, halting further growth.
To quantify the relative amplification of wildtype and mutant strands, we used sequence-specific qPCR with forward primers that selectively bind either the wildtype or mutant sequence, combined with a shared reverse primer (Fig. 2C, Supp. Sec. IV). Two aliquots from each thermocycled sample were analyzed in parallel using both primer sets, allowing independent quantification of wildtype and mutant strand concentrations (Fig. 2D–E). Ct values were determined from the qPCR amplification curves and converted into absolute concentrations using standard curves generated from known DNA concentrations included in the same qPCR run. Control experiments confirmed that the qPCR assay reliably quantifies each strand with high specificity and within the relevant concentration range (Supp. Fig. 2).
qPCR analysis revealed that the wildtype strand was consistently amplified more than the mutant strand over 30 cycles of VCG extension, supporting the hypothesis that competitive strand interactions favor wildtype proliferation and effectively suppress the mutant (Fig. 2E, Supp. Sec. VII).
@@
Figure 3: Mutant suppression depends on both the degree of ‘virtualness’ of the virtual circular genome (VCG) and the proximity of the mutation to 3 end of the oligonucleotide. (A, B) Effect of genome virtualness $V$. Initial VCG pools were constructed from 3, 6, or 12 overlapping oligos in each direction, representing increasing levels of virtualness. (A) Amplification of wt and mutant oligonucleotides as a function of thermal cycle at each level of virtualness, shown as fold-change relative to the initial concentration (y-axis) across thermal cycles (x-axis). Concentrations were inferred from qPCR Ct values, as in Fig. 2E. The initial concentration of mutants was set at 5% of the VCG oligos. (B) wt advantage as a function of virtualness $V$. wt advantage is the ratio of wt to mutant amplification at cycle 30. Different lines represent results from different initial concentrations of mutants. wt advantage is strongest in highly virtual VCGs. In contrast, wt advantage is lost when virtualness is low, approximating a physical circular genome. (C, D) Effect of mutation position. Oligos contain a single mutation at varying positions along the sequence. (C) Amplification of wildtype (wt) and mutant oligos, (D) Amplification advantage of wt over mutant (wt advantage) plotted against the mutation position on the mutant. wt advantage is strongest when the mutation is located at the 3 end.
@@
To investigate the role of VCG architecture in mutation suppression, we varied the system’s ‘virtualness’—defined as the number of overlapping strands encoding the genome—by constructing VCGs with 12, 6, or 3 strand pairs (Fig. 3C, Supp. Table II). All of these sets of strands encode the same 60 bp genome but differ in the number of distinct potential reverse complement binding partners available to each strand. Lower virtualness (e.g., the 3-strand VCG) more closely resembles a real physical circular genome in that it has low redundancy measured by how many different strands cover a given subsequence of the genome. We observed stronger suppression of mutant amplification in higher-virtualness VCGs across all mutant input levels we tested, including 2.5% (Fig. 3A–B), 5% (Supp. Fig. S6), and 25% (Supp. Fig. S7). In the 12-strand VCG, the wildtype outcompeted the mutant by over 10-fold, while in the 3-strand VCG, the amplification advantage for the wildtype was nearly absent (Fig. 3D).
@@
We next examined how the position of the 4 bp mutation within the mutant strand affects its suppression. Using additional mutant variants with mutations placed near the 3 or 5 ends, we found that mutations near the 3 end were most strongly suppressed during extension (Fig. 3A–B). In contrast, the 5-end mutant exhibited amplification levels nearly indistinguishable from the wildtype. This trend held consistently across all VCG virtualness levels and initial mutant proportions (Fig. 3D, Supp. Fig. 5, 6, 7), suggesting that suppression is more effective when the mutation is closer to the 3 end—the site where strand extension initiates. This positional bias aligns with the expectation that errors or mismatches near the 3 end would exhibit a decreased likelihood of propagation.
@@
These results were consistent across all initial mutant concentrations tested (2.5%, 5%, and 25%; Supp. Fig. 5, 6, 7). Together, these findings demonstrate that both VCG virtualness and mutation position govern selective replication within this strand-extension-based system. The results suggest that VCG architectures can impose intrinsic fidelity constraints, naturally suppressing the propagation of mutant alleles relative to the wildtype allele during thermal cycles.
@@ II.1 Simplified VCG Model
Figure 4: Cooperative effects in a VCG simulation reproduce differential suppression of mutant alleles. (A) Our (differential equations-based) simplified VCG model consists of thermal cycles with the following events: annealing, extension, and melting. Annealing: oligos with common sequence overlap $o$ above a minimum length ($o_{min}=2$) can bind to each other to become duplexes. Extension: oligos with 3 ends annealed to their duplex templates extend as primers off of each other to form blunt ends. Melting: Duplexes with fewer than a maximum number ($o_{max}=55$) of paired bases are allowed to melt. Initial oligo pools are chosen to match experimental oligo pools. See Supp. Sec. VIII for further details. (B) (top) Distribution of oligo lengths after the melting step of each of 10 simulated thermal cycles. Relative concentration across each column is normalized by max binned concentration value. (bottom) Concentration of single-stranded ($s$) oligos, concentration of unmelted duplexes, and total concentration of all oligos after the melting step of each thermal cycle. ‘pool stasis’ at cycle 10 indicates all oligos are bound in unmeltable long duplexes ($o\geq o_{max}$). (C) (left) mut and wt region amplification as a function of thermal cycle for three different values of VCG virtualness $V$, defined as number of distinct consensus sequences that make up initial oligo pool. Region amplification is defined as the instantaneous concentration of an allele plus a flanking region (set by the oligo the mut allele initially appears on), normalized by region’s initial concentration. (right) wt advantage (wt region amplification divided by mut region amplification at the end of 10 cycles) as a function of $V$. mut allele introduced in the middle of a single oligo for all $V$ conditions. (D) (left) mut and wt region amplification for three different initial mut positions (5-proximal to 3-proximal) as a function of thermal cycle. (right) wt advantage as a function of initial mut position. $V=12$ across different mut position conditions.
To understand the source of position- and virtualness-dependent mutation fates in the VCG, we build a simplified model of VCG replication (Fig. 4A). In this model, we consider an initial pool of oligos which grows through repeated thermal cycles. This initial oligo pool is based on the experimental oligo pools studied in Sec. I. It contains 12 oligo sequences of length 25 bp which form a consensus VCG sequence of length 60 bp with an offset of 5 bp between each consecutive oligo. In addition, the pool also contains all the reverse complements of the original 12 oligos, plus a small concentration of an additional oligo that is identical to the underlying VCG sequence except at a single position, where it contains a mutant allele.
@@
    First, oligos anneal with any possible complementary partner to form duplexes, following irreversible second-order reaction kinetics until no further annealing reactions are possible. Only oligos which have exact contiguous overlaps above a threshold $o_{min}$ amount are allowed to bind.
@@
    Second, if an oligo has an annealed 3 end, it acts as a primer and is extended using its annealed duplex partner as a template. This extension continues without any errors until the extending oligo reaches either the 5 end of its template or a maximal length $l_{max}$ equal to the length of the consensus VCG sequence (60 bp).
@@
    Finally, during the melting stage, duplexes dissociate into single-stranded oligonucleotides unless their overlap exceeds a specified threshold length $o_{max}$.
In most simulations, we take $o_{min}=2$ and $o_{max}=55$, meaning that oligos must overlap by at least 2 bp in order to anneal and duplexes with more than 55 bp overlap do not melt (as approximately observed in our experimental conditions). We note that annealing in any single thermal cycle is random in that the annealed duplexes are kinetically determined and thus only a small fraction of an oligo may be bound to its perfect complementary pair. For instance, in the initial oligo pool with where oligos are of length 25 bp and are separated by offsets of 5 bp, then any oligo has $9$ potential binding partners, only one of which is its perfect complementary pair. Furthermore, if all oligos are initially present at equal concentration, then only $1/9$ of annealed duplexes will be bound in perfect complementary pairs which cannot extend any further.
Thus, because this system anneals through irreversible kinetics, not all pairs bind with perfect complementarity and hence extension can occur through the formation of duplexes with overhangs. The formation of duplexes with perfect complementarity is further disrupted by the melting phase of thermal cycles on a timescale $\tau_{cycle}$. For more details on how our model is defined and implemented, please see Supp. Sec. VIII.

### v2

@@ Abstract
Modern life largely transmits genetic information from mother to daughter through the duplication of single physically intact molecules that encode information. However, copying an extended molecule requires complex copying machinery and high fidelity that scales with the genome size to avoid the error catastrophe. Here, we explore these fidelity requirements in an alternative architecture, the virtual circular genome, in which no one physical molecule encodes the full genetic information. Instead, information is encoded and transmitted in a collective of overlapping and interacting segments. Using a model experimental system of a complex mixture of DNA oligomers that can partly anneal and extend off each other, we find that mutant oligomers are suppressed relative to a model without collective encoding. Through simulations and theory, we show that this suppression of mutants can be explained by competition for productive binding partners. As a consequence, information can be propagated robustly in a virtual circular genome even at mutation rates expected under prebiotic conditions.
Faithful copying of heritable information is a basic requirement for genomes. Standard accounts of fidelity emphasize enzyme-based mechanisms ranging from nucleotide selectivity[1] and exonucleolytic proofreading[2] to post-replicative mismatch repair[3] acting on a single, continuous template. These mechanisms are powerful but in extant biology are based on sophisticated protein machinery, and they must operate below the well-known error-catastrophe threshold to maintain information[4]. A complementary approach, often overlooked, involves changing the architecture of the genome, i.e., how genetic information is physically laid out. Biology offers many precedents for diverse architectures: the fragmented chromosomes and plasmids of Borrelia burgdorferi[5]; the linear, circular and branched forms of mitochondrial genomes[6]; the partitioning into a micronucleus and macronucleus in ciliates[7]; and the interlocked kinetoplast DNA networks of many parasites[8]. Yet the impact of such architectural choices on how errors arise and propagate remains poorly understood.
@@
Figure 1: Cooperative replication of genetic information in a virtual circular genome. (A) (i) In physically circular genomes found in extant life, genetic information is encoded in long nucleic acid polymers (ii) that are replicated by the extension of short primers. (iii) As each primer is extended to cover the entire genome, a neutral mutant allele in one part of the genome is replicated just as often as its wildtype counterpart. (B) (i) In contrast, the proposed virtual circular genome of a protocell is the consensus sequence (purple) of many short oligomers (black and grey). (ii) Each oligomer (oligo) may act as both primer and template during replication. (iii) Here, we show that the resulting co-operative effects, in which wildtype and mutant oligos compete as both primers and templates, suppress the replication of isolated mutant alleles in favor of wildtype alleles already coded on many oligos.
Alternative architectures for genomic information are also motivated by the search for minimal self-replicating systems, whether envisioned for synthetic biology or as models of early life. Copying an extended template end-to-end typically requires elaborate machinery that can unwind and manage long duplex regions; in modern cells this role is played by the multi-enzyme replisome (helicases, polymerases, gyrases and many other accessory factors). Consequently, even engineered minimal cells currently rely on sizable gene sets to support these processes[9]. Such machinery is unavailable in early evolutionary settings and in more minimal synthetic constructs. Hence it is natural to consider distributed architectures in which information is carried by many short genomic fragments and replication proceeds via many short, parallel extensions rather than a single long uninterrupted pass[10, 11, 12, 13, 14, 15, 16, 17, 18]. We refer to this class of architectures where a population of short, overlapping strands jointly stores and transmits a sequence as a collective encoding.
These considerations motivate the central question we address here: Do collectively encoded genomes, by virtue of their architecture alone, improve replication fidelity? We address this question in the context of the virtual circular genome (VCG) framework[19, 20], in which overlapping genomic fragments map to a circular consensus sequence, even though no single strand contains the whole. We use an experimental DNA-based implementation of a VCG, together with simulations and theory, to compare how wildtype and mutant sequences propagate in a collectively encoded pool.
Our results reveal an inherent asymmetry created by the architecture. Wildtype sequences that match the consensus are distributed across many different strands and hence retain many routes to continued copying. In contrast, mutations appear at singular locations and hence can have reduced routes to copying. The deficit is most severe for changes near the growing 3 end, where the location both cuts off continuation options and increases the likelihood of stalling—premature halts that maroon mutated pieces in short, non-contributing products—consistent with prior observations and models of stall-mediated error suppression and its possible role as a precursor to proofreading[21, 22, 23, 24]. Consequently, mutants are intrinsically suppressed relative to wildtype, with the strongest effects for 3-terminal changes.
Our work shows that by distributing information across interacting oligomers, collective encoding provides an architecture-level route to fidelity that complements enzymatic error correction. This alternative is especially pertinent to early life forms, where elaborate enzymes may be absent, and to synthetic self-replication, where robust information transmission from simple parts is a design goal[25, 26, 27].
@@ I Experimental model
[…]
@@
Figure 3: Mutant suppression depends on both the degree of ‘virtualness’ of the virtual circular genome (VCG) and the proximity of the mutation to 3 end of the oligomer. (A, B) Effect of genome virtualness $V$. Initial VCG pools were constructed from 3, 6, or 12 overlapping oligos in each direction (plus their reverse complements), representing increasing levels of virtualness. (A) Amplification of wt and mutant oligos as a function of thermal cycle at each level of virtualness, shown as fold-change relative to the initial concentration (y-axis) across thermal cycles (x-axis). Concentrations were inferred from qPCR Ct values, as in Fig. 2E. The initial concentration of mutants was set at 5% of the VCG oligos. (B) wt advantage as a function of virtualness $V$. wt advantage is the ratio of wt to mutant amplification at cycle 30. Different lines represent results from different initial concentrations of mutants. wt advantage is strongest in highly virtual VCGs. In contrast, wt advantage is lost when virtualness is low, approximating a physical circular genome. (C, D) Effect of mutation position. Oligos contain a mutant region at varying positions along the sequence. (C) Amplification of wildtype (wt) and mutant oligos. (D) Amplification advantage of wt over mutant (wt advantage) plotted against the mutation position on the mutant. wt advantage is strongest when the mutation is located at the 3 end.
We designed a DNA-based system composed of short, overlapping oligomers (oligos) that form a redundant circular architecture. The full VCG comprises 12 distinct 25 nt oligos and their reverse complements. Each of the 12 oligos in one direction overlaps with its downstream neighbor (i.e., the next oligo in the 5–3 direction) by 20 bases, encoding a 60 bp circular genome (Fig. 2A, Supp. Table I). Thus, each oligo initially has four binding partners that productively allow it to be extended or part of it to be copied, i.e., reverse-complementary oligos that can hybridize and serve as templates and/or as primers, with partial complementarity ranging from 5 to 20 bp. Each oligo also has an additional five binding partners which anneal in unproductive configurations that do not allow for primer extension. Less structured DNA experimental models with pools of oligos have been investigated in prior work[14, 16, 28, 29, 12, 15, 13].
To evaluate how mutations propagate within this architecture, we introduced a 25 nt mutant oligo that differs from one VCG oligo by a single 4 bp substitution near the center of the sequence (Fig. 2B, Supp. Sec. I). Mutant oligos were spiked into VCG mixtures at defined low proportions (2.5%, 5%, or 25% relative to the wildtype) to mimic rare variant emergence.
Extension of VCG oligos was driven by thermal cycling with Bst DNA Polymerase (Large Fragment, NEB M0275), using 10–30 cycles of alternating denaturation ($80~^{\circ}\mathrm{C}$) and annealing/extension ($35~^{\circ}\mathrm{C}$), followed by a final enzyme deactivation step ($90~^{\circ}\mathrm{C}$) for 10 minutes. Notably, no new oligomers are supplied to this system; extension is initiated by overlap-driven annealing of partially complementary VCG oligos (Supp. Sec. III). Gel electrophoresis confirmed that initial 25 nt VCG oligos were extended incrementally over cycles, with products reaching a length of $\sim$45 nt by cycle 10, and further reaching $\sim$60 nt by cycle 20, with gel profiles remaining unchanged through cycle 30 (Fig. 2D). Duplexes up to 45 bp ($T_{\mathrm{m}}=80~^{\circ}\mathrm{C}$) likely remain partially meltable during cycling, allowing oligo reshuffling and continued extension, whereas  60 bp duplexes ($T_{\mathrm{m}}=82~^{\circ}\mathrm{C}$) are too stable to denature, halting further growth. We refer to this final state as ‘pool stasis’.
To quantify the relative amplification of wildtype and mutant alleles, we used sequence-specific qPCR with forward primers that selectively bind either the wildtype or mutant sequence, combined with a shared reverse primer (Fig. 2C, Supp. Sec. IV). Two aliquots from each thermocycled sample were analyzed in parallel using both primer sets, allowing independent quantification of wildtype and mutant allele concentrations (Fig. 2D–E). Ct values were determined from the qPCR amplification curves and converted into absolute concentrations using standard curves generated from known DNA concentrations included in the same qPCR run. Control experiments confirmed that the qPCR assay reliably quantifies each allele sequence with high specificity and within the relevant concentration range (Supp. Fig. S1).
qPCR analysis revealed that the wildtype allele was consistently amplified more than the mutant allele over 30 cycles of VCG extension; competitive oligo interactions favor wildtype proliferation and effectively suppress the mutant (Fig. 2E, Supp. Sec. VI). Notably, this suppression effect is time-dependent: before cycle 10 the mutant is strongly suppressed, while from cycles 10 to 20, wildtype and mutant grow with equal speed until pool stasis. We define a ‘wt advantage’ metric as the ratio of wildtype and mutant amplification at pool stasis (i.e., at cycle 30).
@@
To investigate the role of VCG architecture in mutation suppression, we varied the system’s ‘virtualness’ $V$ – defined as the number of overlapping oligo pairs encoding the genome – by constructing VCGs with 12, 6, or 3 oligo pairs (Fig. 3A, Supp. Table II). All of these sets of oligos encode the same 60 bp genome but differ in the number of distinct potential reverse complement binding partners available to each oligo. Lower virtualness (e.g., the 3-oligo VCG) more closely resembles a real physical circular genome in that it has low redundancy measured by how many different oligos cover a given subsequence of the genome. We observed stronger suppression of mutant amplification in higher-virtualness VCGs across all mutant input levels we tested, including 2.5% (Supp. Fig. S4), 5% (Fig. 3A; Supp. Fig. S5), and 25% (Supp. Fig. S6). In the 12-oligo VCG, the wildtype outcompeted the mutant by over 10-fold, while in the 3-oligo VCG, the amplification advantage for the wildtype was nearly absent (Fig. 3B).
@@
We next examined how the position of the 4 bp mutation within the mutant oligo affects its suppression. Using additional mutant variants with mutations placed near the 3 or 5 ends, we found that mutations near the 3 end were most strongly suppressed during extension (Fig. 3C–D). In contrast, the 5 end mutant exhibited amplification levels nearly indistinguishable from the wildtype. This trend held consistently across all VCG virtualness levels and initial mutant proportions (Fig. 3D, Supp. Fig. S4–6), suggesting that suppression is more effective when the mutation is closer to the 3 end—the site where oligo extension initiates. This positional bias aligns with the expectation that errors or mismatches near the 3 end would exhibit a decreased likelihood of propagation.
@@
Together, these findings demonstrate that both VCG virtualness and mutation position govern selective replication within this oligo-extension-based system. The results suggest that VCG architectures can impose intrinsic fidelity constraints, naturally suppressing the propagation of mutant alleles relative to the wildtype allele during thermal cycles.
@@ II.1 Simplified VCG Model
Figure 4: Cooperative effects in a VCG simulation reproduce differential suppression of mutant alleles. (A) Our (differential equations-based) simplified VCG model consists of thermal cycles with the following events: annealing, extension, and melting. Annealing: oligos with common sequence overlap $o$ above a minimum length ($o_{min}=2$) can bind to each other to become duplexes. Extension: oligos with 3 ends annealed to their duplex templates extend as primers off of each other to form blunt ends. Melting: Duplexes with fewer than a maximum number ($o_{max}=55$) of paired bases are allowed to melt. Initial oligo pools are chosen to match experimental oligo pools. See Supp. Sec. VII for further details. (B) (top) Distribution of oligo lengths after the melting step of each of 10 simulated thermal cycles. Relative concentration across each column is normalized by max binned concentration value. (bottom) Concentration of single-stranded ($s$) oligos, concentration of unmelted duplexes, and total concentration of all oligos after the melting step of each thermal cycle. ‘pool stasis’ in simulations at cycle 10 indicates all oligos are bound in unmeltable long duplexes ($o\geq o_{max}$). (C) (left) mut and wt region amplification as a function of thermal cycle for three different values of VCG virtualness $V$, defined as number of distinct oligo pairs that make up initial oligo pool (either 3, 6, or 12). Region amplification is defined as the instantaneous concentration of an allele plus a flanking region (set by the oligo the mut allele initially appears on), normalized by region’s initial concentration. (right) wt advantage (wt region amplification divided by mut region amplification at pool stasis, here 10 cycles) as a function of $V$. mut allele introduced in the middle of a single oligo for all $V$ conditions. (D) (left) mut and wt region amplification for three different initial mut positions (5-proximal to 3-proximal) as a function of thermal cycle. (right) wt advantage as a function of initial mut position. $V=12$ across different mut position conditions.
To understand the source of position- and virtualness-dependent mutation fates in the VCG, we build a simplified model of VCG replication (Fig. 4A). In this model, we consider an initial pool of oligos which grows through repeated thermal cycles. This initial oligo pool is based on the experimental oligo pools studied in the previous section. It contains 12 oligo sequences of length 25 nt which form a consensus VCG sequence of length 60 bp with an offset of 5 bp between each consecutive oligo. In addition, the pool also contains all the reverse complements of the original 12 oligos, plus a small concentration of an additional oligo that is identical to the underlying VCG sequence except at a single position, where it contains a mutant allele. This is in contrast to the experimental mutant oligo, which differs from its wt counterpart in a 4 nt block.
@@
    First, oligos anneal with any possible complementary partner to form duplexes, following irreversible second-order reaction kinetics until no further annealing reactions are possible. Only oligos which have exact contiguous overlaps above a threshold $o_{min}$ amount are allowed to bind. Here $o_{min}=2$.
@@
    Second, if an oligo has an annealed 3 end, it acts as a primer and is extended using its annealed duplex partner as a template. This extension continues without any errors until the extending oligo reaches either the 5 end of its template or a maximal length $l_{max}$ equal to the length of the consensus VCG sequence (60 nt).
@@
    Finally, during the melting stage, duplexes dissociate into single-stranded oligomers unless their overlap exceeds a specified threshold length $o_{max}$. Here $o_{max}=55$, as approximately observed in experiments.
We note that annealing in any single thermal cycle is random in that the annealed duplexes are kinetically determined. Thus, only a small fraction of an oligo may be bound to its perfect complementary pair. For instance, in the initial oligo pool where oligos are of length 25 nt and are separated by offsets of 5 nt, then any oligo has $9$ potential binding partners, only one of which is its perfect complementary pair. Binding the perfect complementary pair forms a blunt-end duplex which does not allow for further extension. However, each oligo has an additional 8 other binding partners which can bind in equal probability; 4 of those 8 bind in duplexes which allow for extension.
Therefore, because this system anneals through irreversible kinetics, not all pairs bind with perfect complementarity and hence extension can occur through the formation of duplexes with overhangs. The formation of duplexes with perfect complementarity is further disrupted by the melting phase of thermal cycles on a timescale $\tau_{cycle}$. For more details on how our model is defined and implemented, please see Supp. Sec. VII.